OpenStudy (anonymous):

simplify the rational expression, state any restrictions on the variable...t^2+t-6/t^2-9

Nnesha (nnesha):

factor numerator and denominator

Nnesha (nnesha):

\[\huge\rm (a^2 -b^2) = (a+b)(a-b)\] difference of square apply that at the denominator

OpenStudy (anonymous):

t-2/t-3 t=3 t=-3? is what i got for the final

Nnesha (nnesha):

t^2 + t - 6 what are factors of this equation ?

OpenStudy (anonymous):

3

Nnesha (nnesha):

just 3 ?

OpenStudy (anonymous):

i mean you could go 2,3 6

OpenStudy (anonymous):

so my answer up there wasnt right?

Nnesha (nnesha):

3 times -2 = - and 3-2= 1 that would be factor of t^2 + t - 6 :-)

Nnesha (nnesha):

yes that's right

Nnesha (nnesha):

t = 3 t = -3 ??? explain this plz

OpenStudy (anonymous):

So its the right choice? explain what?

OpenStudy (anonymous):

My brother just told me he thinks that was right, he didnt explain

Nnesha (nnesha):

so you don't know how to solve it ?

OpenStudy (anonymous):

No sir, He just said i think its that one, no explaining at all. i would like to know how to though.

Nnesha (nnesha):

t^2 +t -6 find two number if you multiply them you should get -6 and if you add or subtract them you should get middle term which is one

OpenStudy (anonymous):

why is it not 3, -2?

Nnesha (nnesha):

who said it's not ?

OpenStudy (anonymous):

haha, okay, but those two should give us that

Nnesha (nnesha):

it is so bec leading coefficient is 1 just write it as ( t+ 3) and (t-2) this is the numerator \[\frac{ (t+3)(t-2) }{ ? }\] now denominator t^2 - 9 factor it

OpenStudy (anonymous):

ok what do you mean? im sorry.

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

that one gave me t+3 t-3

Nnesha (nnesha):

^^

Nnesha (nnesha):

\[\huge\rm \frac{ (t+3)(t-2) }{ (t+3)(t-3) }\] simplify

OpenStudy (anonymous):

t-2/t-3

Nnesha (nnesha):

ye right now denominator can't be equal to 0 so tha'ts mean \[\huge\rm t-3 \cancel{=} 0\] solve for t

OpenStudy (anonymous):

=3

Nnesha (nnesha):

not equal

OpenStudy (anonymous):

And thats the final answer? :)

OpenStudy (anonymous):

i know. just 3 is the restriction

Nnesha (nnesha):

\[\huge\rm t \cancel{=} 3\] bec if you plug 3 for t 3-3= 0 but we don't like 0 for this question

Nnesha (nnesha):

yes right

Nnesha (nnesha):

gO_Od to go!

OpenStudy (anonymous):

If i open another would you help me? i really appreciate that!! you actually explained. i get that

Nnesha (nnesha):

i'll try :-)